OpenStudy (nikki_skellington):

The holding tanks are congruent in size, and both are in the shape of a cylinder that has been cut in half vertically. The bottom of the tank is a curved surface. What is the volume of both tanks if the radius of tank #1 is 30 feet and the height of tank #2 is 110 feet? You must explain your answer using words, and you must show all work and calculations to receive credit.

OpenStudy (3mar):

Well, I am here.

OpenStudy (nikki_skellington):

:D

OpenStudy (3mar):

so how do you imagine the tanks?

OpenStudy (nikki_skellington):

wait here is the pic

OpenStudy (3mar):

good

OpenStudy (3mar):

so you need to calculate the volume of these two tanks, and we are given the radius and height for both, in addition to that they are half-cylindrical tanks

OpenStudy (nikki_skellington):

how do i do that, like find the volume of the two tanks. Basically meaning is there a formula

OpenStudy (3mar):

in fact, as the two holding tanks are identical, these two halves compose one tank with r=30 and h=110 ok?

OpenStudy (nikki_skellington):

ok

OpenStudy (nikki_skellington):

sorry i just dont know what to do :/

OpenStudy (3mar):

If you know the volume of the cylinder?

OpenStudy (3mar):

What grade are you in?

OpenStudy (nikki_skellington):

heres the thing im in 8th this is for my bro hes in 12th hes struggling with math and he gave up so i had to do it

OpenStudy (nikki_skellington):

thats basically why i barely know anything

OpenStudy (3mar):

You are so kind sister I wish I have one like you Very nice for someone to help you like that

OpenStudy (3mar):

Can I talk to him?

OpenStudy (nikki_skellington):

um let me see what he is doing right now, be right back.

OpenStudy (nikki_skellington):

hes sleeping, as usaul

OpenStudy (3mar):

very cool to sleep or go with my friends and my sister take my responsibilities !!!!!!

OpenStudy (nikki_skellington):

lol, yea i know hes lucky

OpenStudy (3mar):

and you will be a very good mother

OpenStudy (3mar):

anyway

OpenStudy (nikki_skellington):

aww thank you.

OpenStudy (3mar):

you saw the pic I sent?

OpenStudy (nikki_skellington):

yup

OpenStudy (nikki_skellington):

so is the answer 99000

OpenStudy (3mar):

where is \[\pi\]?

OpenStudy (nikki_skellington):

i just dont know how to make the symbol :/

OpenStudy (3mar):

you can just type "pi" or use this icon to show it like this \[\pi\]

OpenStudy (nikki_skellington):

ohhhh ok.

OpenStudy (nikki_skellington):

99000\[\pi\] there

OpenStudy (nikki_skellington):

do you mind helping me with 2 more their short this time?

OpenStudy (3mar):

Do you mind to wake your brother and let him do his homework? ahahahha

OpenStudy (nikki_skellington):

lol

OpenStudy (nikki_skellington):

he'll be mad at me for waking his "beauty" sleep

OpenStudy (3mar):

I am joking! Go ahead! But 1 condition!

OpenStudy (3mar):

Thank you for the medal!